QUADRUPLY BONDED M_{2} COMPLEXES INCORPORATING THIENYLETHYNYL AND THIENYLVINYL CARBOXYLATES

\underline{CARLY R. REED}, MALCOLM H. CHISHOLM, and CLAUDIA TURRO, Ohio State University, Department of Chemistry, Columbus, Ohio 43210.

Quadruply bonded metal complexes, where M = Mo, W, that incorporate thienylethynyl and thienylvinyl carboxylates are being synthesized. Varying the bond order of the C_{2} unit between the thienyl group and the carboxylate tether provides a series of complexes in which a comparative study on the electronic delocalization in the ground and excited states can be conducted. The properties of these complexes will be predicted by DFT calculations and probed by steady-state absorbance, emission, electrochemistry, and ns- and fs-transient absorption.